# Polynomial bounds for surfaces in cusped 3-manifolds

### v2

@@ Abstract.
We find polynomial upper bounds on the number of isotopy classes of connected essential surfaces embedded in many cusped 3-manifolds and their Dehn fillings. Our bounds are universal, in the sense that we obtain the same explicit formula for all 3-manifolds that we consider, with the formula dependent on the Euler characteristic of the surface and similar numerical quantities encoding topology of the ambient 3-manifold. Universal and polynomial bounds have been obtained previously for classical alternating links in the 3-sphere and their Dehn fillings, but only for surfaces that are closed or spanning. Here, we consider much broader classes of 3-manifolds and all topological types of surfaces. The 3-manifolds are called weakly generalized alternating links; they include, for example, many links that are not classically alternating and/or do not lie in the 3-sphere, many virtual links and toroidally alternating links.
@@ 1. Introduction
For decades, the study of essential surfaces in 3-manifolds has led to important topological and geometric consequences. For example, the existence or nonexistence of low genus essential surfaces gives crucial insight into the geometrization of the 3-manifold, due to work of Thurston [38]. Every 3-manifold is now known to have a finite cover containing an embedded essential surface, by significant mathematical achievements involving work of Kahn and Markovic [23], Haglund and Wise [12], and Agol [3], among others.
@@
For fixed Euler characteristic, it is natural to ask how many embedded essential surfaces lie in a given 3-manifold (up to ambient isotopy). For certain 3-manifolds, the number is infinite; for example there are infinitely many nonisotopic Seifert surfaces in a connect sum of knots, as shown in work of Eisner [8]. However, for 3-manifolds with no essential embedded spheres or tori, the number of embedded surfaces with fixed topological type is finite; see for example Jaco and Oertel [21].
Analogues of the question of how many essential surfaces embed in a given 3-manifold have been addressed by others. For example, Masters [26] and then Kahn and Markovic [22] found exponential bounds on the number of immersed essential surfaces. Recently, Dunfield, Garoufalidis, and Rubinstein found quasi-polynomial bounds depending on the 3-manifold, with some restrictions [7]. Restricting to important families of 3-manifolds allows even stronger results. For example, Hass, Thompson and Tsvietkova found polynomial bounds on essential surfaces in the complement of alternating links in the 3-sphere [15, 16]. One dimension lower, a similar question was addressed by Mirzakhani for curves on surfaces [32].
This paper finds polynomial upper bounds on the number of embedded essential surfaces for broad classes of 3-manifolds with torus boundary, and many of their Dehn fillings. These results are novel on several levels. For example, our work immediately applies to the complement of alternating knots in the 3-sphere, and addresses cases not included in work of Hass, Thompson, and Tsvietkova. In particular, the previous work in [15, 16] bounds the number of closed orientable surfaces and spanning surfaces, meaning the boundary is required to be a longitude. Here, we give bounds on any embedded surface, with or without boundary, with no restrictions on boundary slope. Our results also apply to orientable and non-orientable surfaces. This completes the picture for all surfaces within classical alternating knot and link complements.
This paper can also be seen as the next step forward in the same direction as the work of Kahn and Markovic [22] and the work of Dunfield, Garoufalidis, and Rubinstein [7]. The previous step was showing that in addition to exponential bounds for immersed connected surfaces in all hyperbolic 3-manifolds [22], quasi-polynomial bounds for embedded connected surfaces are possible in broad classes of 3-manifolds [7]. (More precisely, the latter work [7] gives an algorithm producing quasi-polynomial count for disconnected surfaces in a fixed 3-manifold, and that count serves as an upper bound for connected surfaces in the given 3-manifold.) Both [22, 7] provide remarkable results, but neither of them explains how the upper bound depends on the 3-manifold; instead, the bounds utilize genus or Euler characteristic of the surface. Hence, once a 3-manifold changes, the expression of the bound changes mysteriously, while the order of dependence on the genus or Euler characteristic of surfaces stays the same. The work [7] conjectures that the dependence on a 3-manifold can be captured through two parameters of the space of measured laminations without boundary in a 3-manifold.
Here, we provide a polynomial upper bound for a broad class for 3-manifolds, which is the natural next step from exponential and quasi-polynomial bounds. More importantly, we achieve the polynomial order because we now know how the bound depends on a 3-manifold. With this, our bound is universal and explicit, i.e. stated as the same straightforward formula for all 3-manifolds that we consider. It depends both on genus or Euler characteristic (exponentially) of surfaces and on two parameters that belong to a 3-manifold (polynomially). This makes our work complementary to what was established in [22, 7]. Note that our two parameters are not relevant to the lamination space, as in the forementioned conjecture: instead, they are related to torus boundary and minimal triangulations of 3-manifold. This is because both our bound and our methods are very different in nature from [22, 7].
Our methods allow us a high level of generality: for example, this is the first paper where the bound applies not only to closed surfaces (as in [7, 15]) or surfaces with one longitudal boundary component, i.e. spanning surfaces (as in [16]), but to all surfaces, with or without boundary, and with any number and type of boundary components. While the bounds for surfaces with several meridianal boundary components are a corollary of bounds for closed surfaces, no work gave bounds for surfaces with multiple boundary components, all with different slopes, before this paper. Our bound also includes not only orientable surfaces (as in [7]), but non-orientable ones as well.
1.1. Main results.
We bound the count of essential surfaces for a broad family of cusped 3-manifolds called weakly generalized alternating link complements, as well as for many closed 3-manifolds that are their Dehn fillings. These first appeared in the work of Howie [18] and Howie and Purcell [19]. They include classes of 3-manifolds that have been considered over several decades, including complements of all classical alternating links and many non-alternating links in the 3-sphere, virtual alternaing links, and others; see below.
Informally, they are complements of links in 3-manifolds such that the link has an alternating diagram on some embedded surface, with mild restrictions. In this paper, we show that for many families of weakly generalized alternating links, if the alternating link diagram has $n$ crossings, then the number of essential surfaces of fixed Euler characteristic $\chi$ is bounded above by an explicit universal polynomial in $n$, of degree at most $-800\chi^{3}+80\chi^{2}$.
The most general statement is:
@@
The factor $X$ is explicitly known for a number of families of cusped 3-manifolds. With this, our work gives simple polynomial bounds depending only on $\chi,g,n$ for:
(1) Alternating links in the 3-sphere. For $L$ a nonsplit prime alternating link, our result extends work of Hass, Thompson, and Tsvietkova, who bound the number of closed surfaces and meridianal surfaces [15], and spanning surfaces with slope a longitude of the knot [16]. Our work gives bounds on arbitrary (connected) essential surfaces with any boundary slope, including both orientable and nonorientable surfaces; see Theorems 6.1 and 6.3. If $Z$ is meridianally incompressible, the number of embeddings is at most $(6n)^{80\chi(Z)^{2}}$; otherwise there is an additional factor in terms of $\chi(Z)$.
(2) Alternating links in a thickened surface $Y$ (i.e. virtual alternating links). Such 3-manifolds were recently investigated by Adams et al [1] and Champanekar, Kofman, and Purcell [6]. Here, we obtain a bound on the number of essential embedded surfaces in Corollary 8.6 below.
@@
(3) Weakly generalized alternating links with a projection surface $\Pi$ that is a Heegaard torus in a lens space, or in a thickened torus, or in a solid torus. The bound is Corollary 8.4. These classes of 3-manifolds appear frequently in low-dimensional topology. For example, alternating knots on Heegaard tori have been studied by Adams [2]. Alternating knots on more general Heegaard surfaces have been studied by Hayashi [17].
As a corollary, we obtain bounds on the number of essential surfaces of fixed Euler characteristic $\chi$ in most closed manifolds obtained by Dehn filling on weakly generalized alternating links, including classical alternating knots; see Theorem 9.1 and Corollary 9.3.
1.2. Comparison to previous work.
As noted above, the work presented here is novel on several levels: It gives polynomial bounds in addition to the existing quasi-polynomial and exponential bounds, applies to closed surfaces as well as those with boundary, applies to both orientable and nonorientable surfaces, and is naturally tailored to working with connected surfaces, which is often more subtle than working with disconnected ones. More importantly, the bound is universal, meaning it is the same formula for all 3-manifolds that we consider, capturing explicit dependence on numerical parameters of surfaces (genus and Euler characteristics) and numerical parameters of 3-manifolds (number of tetrahedra in a minimal triangulation of a chunk, number of crossings in a projection of cusp boundary). Note universal and explicit bounds were also obtained in [15, 16], but not in  [22, 7].
Similar to other work, our upper bounds for connected surfaces are not sharp. Indeed, due to various shortcuts we took to simplify our bounds, and in the combinatorial part of the proof, the final upper bound is likely quite far from sharp. We have not analysed how far from sharp it might be; this would be interesting follow up work.
Finally, the methods we develop here are also new: they have not been used in any of the previous bounds. Some of our techniques can be seen broadly as generalizations of normal surface theory. However the direct application of classical normal surface theory to this problem can only produce universal bounds that are exponential. For example, using triangulations of 3-manifolds and normal surface theory, one can bound the number of isotopy classes of fundamental normal surfaces; see Matveev [27], Jaco and Oertel [21], and Hass, Lagarias, and Pippenger [13]. This can be used to give bounds on the number of all essential surfaces in certain settings, but the bounds are a tower of exponentials in terms of genus. Normal surfaces are also used in the recent work of Dunfield, Garoufalidis, and Rubinstein [7], but, as explained above, the set up is different: instead of a universal bound, normal surfaces allow an algorithm that will count disconnected closed orientable surfaces once a 3-manifold is fixed.
@@
Note that a related, but much simpler problem is bounding the number of incompressible surfaces that can be simultaneously disjointly embedded in a 3-manifold. This problem was originally studied by Kneser, who obtained bounds in 1929 [24]. Here, we bound all essential embeddings, not just disjoint collections.
@@ 1.3. Organisation
In Section 2, we recall the definition of a weakly generalized alternating link, and the decomposition of its complement into chunks introduced by Howie and Purcell [19]. We also review certain techniques for essential surfaces in these link complements: normal form with respect to a chunk decomposition, combinatorial area, and results from Purcell and Tsvietkova [36] giving restrictions on how such surfaces meet chunks. We begin the count of surfaces in Section 4, by bounding the number of subsurfaces making up an essential surface, and the number of labels on the boundaries of such subsurfaces. In Section 5, we restrict the possible number of boundary curves of subsurfaces.
@@ 2.1. Weakly generalized alternating links
Throughout, we work with PL manifolds. We let $Y$ be a compact, orientable, irreducible 3-manifold, possibly with boundary. Embedded in $Y$ is a closed, orientable surface $\Pi$. If $Y$ has boundary, we require $\partial Y$ to be incompressible in $Y-N(\Pi)$, where $N(\cdot)$ denotes an open regular neighborhood. Further, we require $Y-\Pi$ to be irreducible.
A generalized diagram is the projection $\pi\colon L\to N(\Pi)$ of a link $L$ onto the surface $\Pi$ in general position. That is, $L$ can be isotoped through $Y$ to lie in $N(\Pi)$. The image of the projection $\pi(L)$ consists of crossings and arcs between them on $\Pi$.
Observe that the requirement that $Y-\Pi$ be irreducible means that if $\Pi$ is the 2-sphere, then $Y=S^{3}$ and the generalized diagram is actually the standard diagram of a knot in the 3-sphere.
A generalized diagram is alternating if, for each region of $\Pi{\smallsetminus}\pi(L)$, each boundary component of the region is alternating. That is, it can be oriented such that crossings run from under to over in the direction of orientation. An alternating generalized diagram is checkerboard colored if each region of $\Pi{\smallsetminus}\pi(L)$ is oriented so that the induced orientation on the boundary is alternating. See [36, Figure 1] for examples and nonexamples.
@@
Thus it measures the minimal number of intersections of the boundary of any essential compressing disk on one side of $\Pi$. See [36, Example 2.2] for examples.
@@
A generalized diagram $\pi(L)$ on $\Pi$ is defined in [19] to be weakly generalized alternating if
@@
From now on, every link we consider will have such a diagram, i.e. it will be a weakly generalized alternating link. Note that a classical reduced, prime, alternating diagram of a link $L$ on $\Pi=S^{2}$ in $S^{3}$ is an example of a weakly generalized alternating link. As noted in the introduction, there are many more examples, including alternating links on Heegaard tori in $S^{3}$ and in lens spaces. When the diagram is cellular, meaning all regions of $\Pi-\pi(L)$ are disks, these are examples of toroidally alternating knots studied by Adams [2]. Note that weakly generalized alternating knots do not necessarily need to have cellular diagrams. However, the representativity condition means that not every toroidally alternating knot is also weakly generalized alternating. Virtual alternating knots are further examples of weakly generalized alternating knots; these are alternating knots in a thickened surface $\Sigma\times(-1,1)$, projected onto $\Pi=\Sigma\times\{0\}$. They have also received attention recently, for example by Adams [1] and Champanerkar, Kofman, and Purcell [6].

### v3

@@ Abstract.
It is natural to ask how many isotopy classes of embedded essential surfaces lie in a given 3-manifold. The first bounds on the number of such surfaces were exponential, using normal surfaces. More recently, by restricting to alternating link complements in 3-sphere, Hass, Thompson and Tsvietkova obtained polynomial bounds, but for a limited class of surfaces: closed and spanning ones. Here, we complete the picture for classical alternating links and extend these results to other classes of cusped 3-manifolds. We give explicit polynomial bounds on all embedded essential surfaces, closed or any boundary slope, orientable or non-orientable. Our 3-manifolds are complements of links with alternating diagrams on wide classes of surfaces in broad families of 3-manifolds. This includes all alternating links in 3-sphere as well as many non-alternating ones, alternating virtual knots, many toroidally alternating knots, and most Dehn fillings of such manifolds.
@@ 1. Introduction
For decades, the study of essential surfaces in 3-manifolds has led to important topological and geometric consequences. For example, the existence or nonexistence of low genus essential surfaces gives crucial insight into the geometrization of the 3-manifold, due to work of Thurston [52], and every 3-manifold is known to have a finite cover containing an embedded essential surface, by work of Kahn and Markovic [32], Haglund and Wise [21], and Agol [3]. It is natural to ask how many connected embedded essential surfaces lie in a given 3-manifold for fixed Euler characteristic (up to ambient isotopy). For certain 3-manifolds, the number is infinite; for example there are infinitely many nonisotopic Seifert surfaces in a connect sum of knots, as shown in work of Eisner [16]. However, for 3-manifolds with no essential embedded spheres or tori, the number of embedded surfaces with fixed topological type is finite; see for example Jaco and Oertel [30].
Using triangulations of 3-manifolds and normal surface theory, one can bound the number of isotopy classes of fundamental normal surfaces; see Matveev [40], Jaco and Oertel [30], and Hass, Lagarias, and Pippenger [22]. This can be used to give bounds on the number of all essential surfaces in certain settings, but the bounds are a tower of exponentials in terms of genus. Hass, Thompson, and Tsvietkova recently significantly improved upon these results by restricting to alternating link complements. In this setting, they showed that the number of isotopy classes of closed surfaces is polynomial in the number of crossings of an alternating diagram [24]. They then generalized to give similar results for spanning surfaces, i.e. those surfaces with a single boundary component forming a longitude for the knot [25]. However, even in the alternating setting, the most general case remained open, namely counting surfaces of any topological type.
In this paper, we complete the program for alternating links. We find polynomial bounds on the number of isotopy classes of embeddings of any surface, orientable or nonorientable, with or without boundary, and boundary forming any slope on a neighbourhood of the link. We do this by broadening our perspective. Rather than restricting to tools available in the classical alternating setting, we develop and extend more general tools to cusped 3-manifolds that can be described by similar combinatorial patterns as alternating links. The combinatorics required is an alternating diagram on an arbitrary closed surface lying in an arbitrary 3-manifold, with mild hypotheses, namely weakly generalized alternating links defined below. As a consequence, our results immediately extend to give polynomial bounds for broader families of 3-manifolds that have received much recent attention: virtual alternating links, many toroidally alternating links, and most Dehn fillings of these manifolds and classical alternating links.
Our bounds are universal, in the sense that given a fixed topological surface, we obtain the same formula bounding isotopy classes of its embeddings in any of the above 3-manifolds. They are also explicit and effective. The formula is polynomial in the crossing number of a diagram, with the degree of the polynomial and the constants of multiplication depending only on the Euler characteristic of the surface, and most generally, on topology of the ambient 3-manifold.
1.1. Main results
Our first main theorem completes the program started by Hass, Thompson, and Tsvietkova [24, 25]. It gives polynomial bounds for surfaces in alternating link complements, with no restrictions on surface.
Theorem 6.4.
$(6n)^{80\chi(Z)^{2}}2^{-4\chi(Z)+2}$
We also restrict to meridianally incompressible surfaces, obtaining a stronger bound of $(6n)^{80\chi(Z)^{2}}$ in Theorem 6.1. Such surfaces include all quasifuchsian surfaces, giving a bound that applies in a more geometrical setting; see Corollary 6.3. Quasifuchsian surfaces are known to be very prevalent in cusped hyperbolic 3-manifolds by work of Cooper and Futer [14], and in closed hyperbolic 3-manifolds by work of Kahn and Markovic [32].
The methods to prove Theorem 6.4 are much more general than the classical alternating setting. They can be used immediately to give bounds in other important settings, such as that of virtual alternating knots.
Corollary 8.6.
Let $\pi(L)$ be a virtual link $L$, with a weakly prime alternating projection onto a surface $\Pi=F\times\{0\}\subset Y=F\times[-1,1]$ with $n$ crossings. Fix a connected, orientable topological surface $Z$ with Euler characteristic $\chi$. Then up to isotopy, the number of ways to properly embed $(Z,\partial Z)$ as an essential, meridianally incompressible surface in $(Y-N(L),\partial N(L))$ is at most:
$(2(g+1))^{-4\chi}\cdot(6n)^{-800\chi^{3}+80\chi^{2}}$
If in addition $Y-L$ admits a hyperbolic structure, then the above bounds quasifuchsian surfaces.
$(4(g+1))^{-4\chi(Z)+2}(6n)^{-800\chi^{3}+80\chi^{2}}$
Interest in the complements of knots in spaces besides $S^{3}$, with diagrams on surfaces besides $S^{2}$, has been growing in recent years. Some recent results generalize classical knot theory, for example the work of Boden and Karimi [7]. Other research concerns quantum invariants, such as [9, 5], cobordisms, e.g. [53, 11], geometric topology, as in [26, 34, 28], as well as other aspects of these 3-manifolds [6, 12]. One widely studied class of such 3-manifolds is virtual links. They were described by Kauffman in the 1990s as natural generalisations of Gauss codes for classical knot diagrams [35]. They were subsequently shown to be equivalent to diagrams of links in thickened surfaces up to moves called stabilisation and destabilisation [33, 11], and Kuperberg showed that there is a unique minimal genus surface $\Pi$ for which the link embeds in $\Pi\times[-1,1]$ with diagram on $\Pi\times\{0\}$, admitting no destabilisations [37]. In the alternating case, provided the diagram is weakly prime (definition recalled below), such a link satisfies all conditions to ensure it is an instance of a weakly generalized alternating link, defined by Howie and Purcell [27].
Indeed, the tools we develop here apply most broadly in the setting of weakly generalized alternating links. Informally, such links are complements of links in general 3-manifolds that have an alternating diagram on some embedded surface, with mild restrictions involving requirements on connectivity and interaction with compression discs for the projection surface. Our most general result is:
@@ 1.1. Main results
The factor $X$ is explicitly known for a number of families of cusped 3-manifolds. For example, it can be used to give bounds for weakly generalized alternating links with a projection surface $\Pi$ that is a Heegaard torus in a lens space, thickened torus, or solid torus; the bound is Corollary 8.4. These classes of 3-manifolds appear frequently in low-dimensional topology. For example, they include instances of toroidally alternating links introduced by Adams [2].
Again, a direct corollary is a bound on the number of quasifuchsian surfaces in the hyperbolic case; see Corollary 7.4.
Finally, because our bounds apply to all topological surfaces, with any boundary slope on the knot or link complement, we are also able to obtain bounds on the number of essential surfaces in 3-manifolds obtained by Dehn filling such link complements. See Theorem 9.1 and Corollary 9.3.
1.2. Comparison to other results
Analogues of the question of how many connected essential surfaces embed in a given 3-manifold have been addressed by others. Masters [39] and then Kahn and Markovic [31] found exponential bounds on the number of immersed essential surfaces in 3-manifolds that are hyperbolic and closed. More recently, Dunfield, Garoufalidis, and Rubinstein showed that quasi-polynomial upper bounds exist for connected, orientable surfaces for a certain wide class of cusped 3-manifolds that includes hyperbolic knot complements in the 3-sphere [15]. More precisely, they give an algorithm producing a quasi-polynomial count for disconnected surfaces in a fixed 3-manifold. However, all these bounds depend on the 3-manifold. When the 3-manifold changes, the expression of the bound changes in a way that is not well-understood, while the order of dependence on the genus or Euler characteristic stays the same. In contrast, our bounds are explicit and universal, depending on crossing number, and, most generally, the bound on surfaces $X$ from the ambient 3-manifold. We achieve this through developing techniques that are completely different from those of  [31] and  [15].
The methods we develop here are also new: they have not been used in any of the previous bounds. Some of our techniques can be seen broadly as generalizations of normal surface theory, though we do not triangulate the 3-manifold, looking instead at a more general decomposition. Moreover, the direct application of classical normal surface theory to this problem does not yield polynomial bounds, as explained above.
One dimension lower, a similar question was addressed by Mirzakhani for curves on surfaces [45], with numerous extensions, for example [51, 17].
@@ 1.2. Comparison to other results
A related, but simpler problem is bounding the number of incompressible surfaces that can be simultaneously disjointly embedded in a 3-manifold. It was originally studied by Kneser, who obtained bounds in 1929 [36]. Here, we bound all essential embeddings, not just disjoint collections.
1.3. How good are these bounds?
The upper bounds of this paper are likely very far from sharp: our goal is to establish polynomial bounds rather than obtain the most sharp bounds, which would require very delicate analysis at certain stages of the proof. Rather than pursue this analysis, we take frequent shortcuts that preserve the polynomial order of the bound, but at the cost of growing constants and exponents. However, we note that this is still state of the art: there are no known sharp bounds for connected surfaces in any setting, and we have not come across even open conjectures on sharpness in the literature. What we do provide are explicit, effective, computable, and polynomial bounds in the classical, virtual, and toroidally alternating settings. This is novel. Many of the other important results above only prove the existence of certain constants without deriving formulae, where our bounds are immediately computable.
@@ 1.4. Organisation
In Section 2, we recall the definition of a weakly generalized alternating link, and the decomposition of its complement into chunks introduced by Howie and Purcell [27]. We also review certain techniques for essential surfaces in these link complements: normal form with respect to a chunk decomposition, combinatorial area, and results from Purcell and Tsvietkova [49] giving restrictions on how such surfaces meet chunks. We begin the count of surfaces in Section 4, bounding the number of subsurfaces making up an essential surface, and the number of labels on the boundaries of subsurfaces. In Section 5, we restrict the possible number of boundary curves of subsurfaces.
@@ 2.1. Weakly generalized alternating links
Throughout, we work with PL manifolds. Let $Y$ be a compact, orientable, irreducible 3-manifold, possibly with boundary. Embedded in $Y$ is a closed, orientable surface $\Pi$. If $Y$ has boundary, we require $\partial Y$ to be incompressible in $Y-N(\Pi)$, where $N(\cdot)$ denotes an open regular neighborhood. We also require $Y-\Pi$ to be irreducible.
A generalized diagram is the projection $\pi\colon L\to N(\Pi)$ of a link $L$ onto the surface $\Pi$ in general position. That is, $L$ can be isotoped through $Y$ to lie in $N(\Pi)$. The image of the projection $\pi(L)$ consists of crossings and arcs between them on $\Pi$. Observe that the requirement that $Y-\Pi$ be irreducible means that if $\Pi$ is the 2-sphere, then $Y=S^{3}$ and the generalized diagram is actually the standard diagram of a knot in the 3-sphere.
A generalized diagram is alternating if, for each region of $\Pi{\smallsetminus}\pi(L)$, each boundary component of the region is alternating. That is, it can be oriented such that crossings run from under to over in the direction of orientation. An alternating generalized diagram is checkerboard colored if each region of $\Pi{\smallsetminus}\pi(L)$ is oriented so that the induced orientation on the boundary is alternating. See [49, Figure 1] for examples and nonexamples.
@@
Thus it measures the minimal number of intersections of the boundary of any essential compressing disk on one side of $\Pi$. See [49, Example 2.2] for examples.
@@
A generalized diagram $\pi(L)$ on $\Pi$ is weakly generalized alternating [27] if
@@
From now on, every link we consider will be weakly generalized alternating. A classical reduced, prime, alternating diagram of a link $L$ on $\Pi=S^{2}$ in $S^{3}$ is an example of a weakly generalized alternating link. As noted in the introduction, there are many more examples, including alternating links on Heegaard tori in $S^{3}$ and in lens spaces. When the diagram is cellular, meaning all regions of $\Pi-\pi(L)$ are disks, these are examples of toroidally alternating knots [2]. Virtual alternating knots are further examples of weakly generalized alternating knots; these are alternating knots in a thickened surface $\Sigma\times(-1,1)$, projected onto $\Pi=\Sigma\times\{0\}$. Their geometry has also received attention recently, for example [1, 13]. We will revisit them again in Section 8.2.